OpenStudy (anonymous):

courtney and jamal are selling boxes of fruit for their school. each box they sell brings in more money to the school. they will each add seperate contributions when they turn in their money to the school. the equation of the line that describes the amount of money courtney will raise if she sells x boxes of fruit is y=5x+40. Suppose the graph of the equation that describes how much money Jamal will raise is parallel to Courtney's line, and if he sells 4 boxes, he will raise $65. What line describes how much money Jamal will raise for selling x boxes?

OpenStudy (anonymous):

Let's see, if Courtney sells 4 boxes, then she would make (y = 5(4) + 40) $60. Hmmm, so if Jamal sells 4 boxes.. let's see if this equation works. y = 5(4) + 45 y = 20 + 45 y = 65 So Jamal earns y = 5x + 45 (x = number of boxes sold, y = total money earned), I think.